find the simple interest rate if a loan of £6100 gains interest of £1098 in 3 years
Answer by stanbon(75887)

find the simple interest rate if a loan of £6100 gains interest of £1098 in 3 years
=======
Interest = P*r*t
1098 = 6100*r*3
----
r = 1098/(3*6100) = 6%
---------------------
Cheers,
Stan H.
----------------------